OpenStudy (anonymous):

How to graph y=3cos3x-2. I'm having some difficulties finding the intercepts

OpenStudy (anonymous):

Amplitude = 3 Period = 2pi/3 2 units moving down

OpenStudy (anonymous):

What shall I do next? @ganeshie8

OpenStudy (anonymous):

Should I count by 1/4?

ganeshie8 (ganeshie8):

to find x-intercepts put y=0

ganeshie8 (ganeshie8):

3cos3x-2 = 0 cos3x = 2/3

ganeshie8 (ganeshie8):

you wanto solve this in (0, 2pi) ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

put 3x=t, cos t = 2/3

ganeshie8 (ganeshie8):

solve t in (0, 6pi) and divide each solution by 3

OpenStudy (anonymous):

Could you please further explain this bit

OpenStudy (anonymous):

Pi,2pi,3pi, .....6pi?

OpenStudy (anonymous):

And I divide each by 3?

ganeshie8 (ganeshie8):

sure :) #First solution in \(t\): \(\large \cos t = \frac{2}{3}\) \(\large t = \arccos(\frac{2}{3}) = 0.841\)

ganeshie8 (ganeshie8):

lets first find all possible solutions in \(t\), after that we can divide them by 3 to get solutions in \(x\) okay ?

ganeshie8 (ganeshie8):

see if the first solution makes sense ? :)

OpenStudy (anonymous):

Yes that will be 3x=0.84

OpenStudy (anonymous):

So our first intercept will be x= 0.28

ganeshie8 (ganeshie8):

yup ! you got it ! but thats just one intercept. we need to find all other \(t\) values in (0, 6pi)

OpenStudy (anonymous):

Sweet :)

ganeshie8 (ganeshie8):

use below to find two other \(t\) values : \(\large \cos (t) = \cos(t + 2\pi)\)

OpenStudy (anonymous):

So my next t will be ?

OpenStudy (anonymous):

I go by 2pi?

ganeshie8 (ganeshie8):

yes for second solution : add 2pi to the first solution,

OpenStudy (anonymous):

If so then it will be x= 2pi/3

ganeshie8 (ganeshie8):

for third solution : add 2pi to the second solution (which is same as adding 4pi to the first solution)

ganeshie8 (ganeshie8):

nope, u need to add 2pi to the first SOLUTION..

OpenStudy (anonymous):

Ok. So I should go pi, 2pi, 3pi,etc

OpenStudy (anonymous):

I'm confused

ganeshie8 (ganeshie8):

#First solution in \(t\) : \(\large \cos t = \frac{2}{3}\) \(\large t = \arccos(\frac{2}{3}) = 0.841\) Second solution in \(t\) : \(\large 0.841 + 2\pi \) Third solution in \(t\) : \(\large 0.841 + 2\pi + 2\pi\)

ganeshie8 (ganeshie8):

All the 3 solutions above ^^

OpenStudy (anonymous):

Alright now I seem to get it

ganeshie8 (ganeshie8):

there will be 3 more, but before that, see if the previous ones make sense :)

OpenStudy (anonymous):

How do we know when to go by 2pi? Why can't we use pi?

ganeshie8 (ganeshie8):

very good question :) \(2\pi\) is the period of \(\cos t\) function, so, its value REPEATS every \(2\pi\) units

OpenStudy (anonymous):

I thought the period was 2pi/3

ganeshie8 (ganeshie8):

2pi/3 is the period of cos(3x) 2pi is the period of cos(t)

ganeshie8 (ganeshie8):

we had let t = 3x, to make the calculations easy

OpenStudy (anonymous):

Ill try solve this tomorrow in the morning and will send you my answer :)

OpenStudy (anonymous):

Thanks for your help. You are always amazing

ganeshie8 (ganeshie8):

sure.. take ur time ! good luck :)

ganeshie8 (ganeshie8):

no, you're amazing xD

OpenStudy (anonymous):

Hi @ganeshie8

OpenStudy (anonymous):

I tried but I just couldn't get it

OpenStudy (anonymous):

Let us try graph y=sinx + 1

OpenStudy (anonymous):

@ganeshie8